# Speed, Accuracy, and Complexity

### v2

@@ 4.  Ability and Effort
Second, we wish to speak to the ambiguity in the literature regarding the relation between ability and response time. In many contexts slower response times are taken to be indicative of greater ability or sophistication. For instance, more financially literate individuals will by definition be at a better position for understanding information about alternative financial assets (e.g. Darriet et al., 2020). Also in strategic settings, it is often found that individuals who commit more time to a problem tend to be more successful or sophisticated (e.g. Rubinstein, 2016; Agranov et al., 2015; Alós-Ferrer and Bruckenmaier, 2021; Esteban-Casanelles and Gonçalves, 2022; Gill and Prowse, 2023).
@@
We note that Proposition 2 is consistent with recent research in psychology that highlights the importance of taking into account task complexity when interpreting the relation between ability and response time — see e.g. Dodonova and Dodonov (2013), Goldhammer et al. (2014), Kyllonen and Zu (2016), and Junghaenel et al. (2023). In short, these studies find that, in simple tasks, faster responses are indicative of greater ability, but in complex tasks, the opposite is true. Our simple model can rationalise this nuanced prediction.
@@ 5.  General Costs of Information Acquisition
Specifically, we consider a decision-maker faces uncertainty about an unknown parameter or event $\theta\in\Theta$ and can acquire a Blackwell experiment $\pi$ to learn about $\theta$, which, following existing literature (e.g. Kamenica and Gentzkow, 2011; Caplin et al., 2022; Lipnowski and Ravid, 2023), we model as a distribution over posterior beliefs, i.e. $\pi\in\Pi:=\Delta(\Delta(\Theta))$ such that $\mathbb{E}_{\pi}[p]=p_{0}$. We denote by $C:\Pi\to\overline{\mathbb{R}}$ the cost of the experiment, which we assume is strictly increasing with respect to the Blackwell order $\geq_{B}$, continuous with respect to the Lévy-Prokhorov metric, strictly convex, and normalised such that a fully uninformative experiment bears zero cost, $C(\delta_{p_{0}})=0$. We model the benefit of information as a function $B:\Pi\to\mathbb{R}$ that is bounded, linear, and strictly increasing with respect to the Blackwell order, which arises as the decision-maker’s expected payoff from choosing an action $a\in A$ in order to maximise their expected payoff given posterior beliefs $p\sim\pi$, i.e. $B(\pi)=\mathbb{E}_{\pi}[\max_{a\in A}\mathbb{E}_{p}[u(a,\theta)]]$.
@@
In order to connect this framework to time, we assume that the decision-maker’s expected stopping time is proportional to the total cost of information, i.e. $T^{*}(\kappa):=\mathbb{E}[\tau^{*}]\propto\kappa C^{*}(\kappa)$. The scale $\kappa$ is then interpreted as a measure of problem complexity, scaling the marginal cost of information. In the case of the Wald problem, this is simply reciprocal of the square of the signal-to-noise ratio, $\kappa=(\mu/\sigma)^{-2}$, which extends to a setting with many states (Morris and Strack, 2019).171717 See Pomatto et al. (2023) for an axiomatic characterisation of such cost function.
@@
Returning to the binary decision problems introduced in Section 2, we suppose $A=\Theta=\{a,b\}$ and $\tilde{p}=1/2$. If one assumes costs of information are posterior-separable (Caplin et al., 2022), i.e. given by $C(\pi)=\mathbb{E}_{\pi}[c(p)]$ for some $c:[0,1]\to\mathbb{R}_{+}$ satisfying $c^{\prime},c^{\prime\prime}>0$, and $c(p_{0})=0$, then $T^{*}(\kappa)$ is single-peaked if and only if $c(p)/c^{\prime}(p)$ is quasiconcave on $[1/2,1]$.
The argument is simple. First, we recall the well-known result that additively separable information costs, it is without loss of optimally to choose information structures with at most $|A|$ signals, in this case $\overline{p}$ and $\underline{p}$. Symmetry of the problem ($\tilde{p}=p_{0}=1/2$) implies at an optimum $\overline{p}=1-\underline{p}$. The observation follows from the first-order condition $02-\kappa c^{\prime}(p)=$ 0.
While this implies that single-peakedness of $T^{*}(\kappa)$ is not generic, the most common posterior-based information cost functions do have that $c(p)/c^{\prime}(p)$ is quasiconcave on $[1/2,1]$, e.g. entropy costs, $c(p)=p\ln(p)+(1-p)\ln(1-p)-\ln(1/2)$, and quadratic or variance costs, $c(p)=(p-p_{0})^{2}$ — see Figure 3.
@@ 6.  Identifying Complexity
Figure 4 clearly illustrates the result: the higher the indifference point $\tilde{p}$, the more likely the decision-maker is to choose $b$. The effect of problem complexity is then to affect how much more the decision-maker is likely to choose $b$ when the indifference point is increased. It is immediate that the more complex the problem is, insofar as the the choice probability is interior, the stronger the dependence of the choices on the indifference point is, and the steeper the lines are. This is captured by the cross-partial derivative $\partial^{2}\mathbb{P}(\alpha_{\tau}=b)/(\partial\tilde{p}\partial(\mu/\sigma))$ being positive. The proof of this fact follows by obtaining a closed-form characterisation of this derivative using implicit differentiation. The important point is that this result is robust to aggregating over heterogenous agents with different preferences, insofar as these are distributed the same across the two problems.
@@ 7.  Final Remarks
We believe that our results have important implications for the use of response times in the social sciences. While response times provide an important source of information that is able to shed light on the cognitive processes underlying decision-making, our results support existing evidence that, on their own, they are not a reliable indicator of problem complexity or ability. Instead, we propose to use small incentive distortions to infer problem complexity. These are easy to implement in real-life settings and could be used to study which features make specific problems more or less complex, and evaluate if particular policies improve them. In current ongoing, Gonçalves et al. (2024) not only corroborate the results in Theorem 1 — already evidenced in the above-referenced literature — but also demonstrate that incentive manipulation can indeed be used to reveal problem complexity.
@@ A.5.  Proof of Proposition 1
Suppose that $V_{a}$ has a kink at $p\in(\underline{p},\overline{p})$, i.e. $V_{a}^{\prime}(p^{+})\neq V_{a}^{\prime}(p^{-})$. Then, by convexity, $V_{a}^{\prime}(p^{+})>V_{a}^{\prime}(p^{-})$, and, for any $k\in(V_{a}^{\prime}(p^{-}),V_{a}^{\prime}(p^{+}))$ and significantly high $M>0$, $V_{a}(p^{\prime})\geq V_{a}(p)+k(p^{\prime}-p)+(M/2)(p^{\prime}-p)^{2}=:\psi(p^{\prime})$ for $p^{\prime}$ in a neighborhood of $p$. Then, we would obtain that $\tilde{c}-(\mu/\sigma)^{2}(p(1-p))^{2}M<0$ for sufficiently high $M$, contradicting the fact that $V_{a}$ is a viscosity supersolution. Hence, $V_{a}$ is $\mathcal{C}^{1}$ except possibly at $\{\underline{p},\overline{p}\}$.
Since $V_{a}^{\prime\prime}$ exists a.e., from Crandall et al. (1992, p. 15), the HJB will hold with equality whenever $V_{a}$ is twice differentiable, i.e. almost everywhere on $(\underline{p},\overline{p})$. Furthermore, $0\leq V_{a}^{\prime\prime}\leq(\mu/\sigma)^{-2}\tilde{c}\max\{(\underline{p}(1-\underline{p}))^{-2},(\overline{p}(1-\overline{p}))^{-2}\}$, and so $V_{a}^{\prime\prime}$ is bounded whenever it exists. Take any sequence $(p_{n})\subset(\underline{p},\overline{p})$ such that $p_{n}\to p^{*}\in(\underline{p},\overline{p})$. Then, there is a subsequence $(p_{m})$ such that $V_{a}^{\prime}(p_{m})\to V_{a,\infty}^{\prime}$, $V_{a}^{\prime\prime}(p_{m})\to V_{a,\infty}^{\prime\prime}$, and, by continuity, $0=-\tilde{c}+(\mu/\sigma)^{2}\left(p^{*}(1-p^{*})\right)^{2}V_{a,\infty}^{\prime\prime}$. Hence $(V_{a,\infty}^{\prime},V_{a,\infty}^{\prime\prime})$ belongs to both the super- and subjet of $V_{a}$ at $p^{*}$, implying that $V_{a}$ is twice continuously differentiable everywhere except possibly at $\{\underline{p},\overline{p}\}$.
@@ A.9.  Proof of Proposition 3
Let $\pi_{I}$ denote a fully informative experiment — i.e. assigning probability $p_{0}(\theta)$ to posterior belief $1_{\theta}$ — and $\pi_{U}$ a fully uninformative experiment, assigning probability 1 to the prior $p_{0}$. Take any $\overline{C}>0$ such that $\exists\overline{\kappa}$ such that $\overline{\kappa}C^{*}(\overline{\kappa})>\overline{C}$. Then, there are $\kappa<\overline{\kappa}$ and $\kappa^{\prime}>\overline{\kappa}$ such that $\kappa C^{*}(\kappa)\vee\kappa^{\prime}C^{*}(\kappa^{\prime})<\overline{C}$. To see this note that, by Berge’s maximum theorem, $\kappa C^{*}(\kappa)$ is continuous in $\kappa$. Taking $\kappa_{n}\downarrow 0$, $V(\kappa_{n})\uparrow V(0)=B(\pi_{I})$, and so $V(0)-V(\kappa_{n})=\kappa_{n}C^{*}(\kappa_{n})\downarrow 0$. Take now $\kappa_{n}\uparrow\infty$. Note that $0\leq\lim_{n\to\infty}V(\kappa_{n})/\kappa_{n}=\lim_{n\to\infty}B(\pi^{*}(\kappa_{n}))/\kappa_{n}-C^{*}(\kappa_{n})=\lim_{n\to\infty}-C^{*}(\kappa_{n})\leq 0$, and so it must be that $\pi^{*}(\kappa_{n})\to\pi_{U}$. Consequently, $\kappa_{n}C^{*}(\kappa_{n})\leq B(\pi^{*}(\kappa_{n}))\to 0$, a contradiction. ∎
@@ Appendix B.  Discounting
Furthermore, since payoffs are symmetric (i.e. $u(a,a)-u(b,a)=u(b,b)-u(a,b)$ and $u(a,b)=u(b,a)$), the stopping thresholds will also be symmetric around $1/2$, i.e. $\underline{p}=1-\overline{p}$. Then,

### v3

@@ 4.  Ability and Effort
Second, we wish to speak to the ambiguity in the literature regarding the relation between ability and response time. In many contexts slower response times are taken to be indicative of greater ability or sophistication. For instance, more financially literate individuals will by definition be at a better position for understanding information about alternative financial assets (e.g., Darriet et al., 2020). Also in strategic settings, it is often found that individuals who commit more time to a problem tend to be more successful or sophisticated (e.g., Rubinstein, 2016; Agranov et al., 2015; Alós-Ferrer and Bruckenmaier, 2021; Esteban-Casanelles and Gonçalves, 2022; Gill and Prowse, 2023).
@@
We note that Proposition 2 is consistent with recent research in psychology that highlights the importance of taking into account task complexity when interpreting the relation between ability and response time — see e.g., Dodonova and Dodonov (2013), Goldhammer et al. (2014), Kyllonen and Zu (2016), and Junghaenel et al. (2023). In short, these studies find that, in simple tasks, faster responses are indicative of greater ability, but in complex tasks, the opposite is true. Our simple model can rationalise this nuanced prediction.
@@ 5.  General Costs of Information Acquisition
Specifically, we consider a decision-maker faces uncertainty about an unknown parameter or event $\theta\in\Theta$ and can acquire a Blackwell experiment $\pi$ to learn about $\theta$, which, following existing literature (e.g., Kamenica and Gentzkow, 2011; Caplin et al., 2022; Lipnowski and Ravid, 2023), we model as a distribution over posterior beliefs, i.e., $\pi\in\Pi:=\Delta(\Delta(\Theta))$ such that $\mathbb{E}_{\pi}[p]=p_{0}$. We denote by $C:\Pi\to\overline{\mathbb{R}}$ the cost of the experiment, which we assume is strictly increasing with respect to the Blackwell order $\geq_{B}$, continuous with respect to the Lévy-Prokhorov metric, strictly convex, and normalised such that a fully uninformative experiment bears zero cost, $C(\delta_{p_{0}})=0$. We model the benefit of information as a function $B:\Pi\to\mathbb{R}$ that is bounded, linear, and strictly increasing with respect to the Blackwell order, which arises as the decision-maker’s expected payoff from choosing an action $a\in A$ in order to maximise their expected payoff given posterior beliefs $p\sim\pi$, i.e., $B(\pi)=\mathbb{E}_{\pi}[\max_{a\in A}\mathbb{E}_{p}[u(a,\theta)]]$.
@@
In order to connect this framework to time, we assume that the decision-maker’s expected stopping time is strictly increasing in the total cost of information, i.e., $T^{*}(\kappa):=\mathbb{E}[\tau^{*}]=g(\kappa C^{*}(\kappa))$. The scale $\kappa$ is then interpreted as a measure of problem complexity, scaling the marginal cost of information. In the case of the Wald problem, this is simply reciprocal of the square of the signal-to-noise ratio, $\kappa=(\mu/\sigma)^{-2}$ — which extends to a setting with many states (Morris and Strack, 2019)171717 See Pomatto et al. (2023) for an axiomatic characterisation of such cost function. — and expected stopping time is linear in the (ex-ante) total costs of information.
@@
Returning to the binary decision problems introduced in Section 2, we suppose $A=\Theta=\{a,b\}$ and $\tilde{p}=1/2$. If one assumes costs of information are posterior-separable (Caplin et al., 2022), i.e., given by $C(\pi)=\mathbb{E}_{\pi}[c(p)]$ for some $c:[0,1]\to\mathbb{R}_{+}$ satisfying $c^{\prime},c^{\prime\prime}>0$, and $c(p_{0})=0$, then $T^{*}(\kappa)$ is single-peaked if and only if $c(p)/c^{\prime}(p)$ is quasiconcave on $[1/2,1]$.
The argument is simple. First, we recall the well-known result that additively separable information costs deliver that it is without loss of optimally to choose information structures with at most $|A|$ signals, in this case $\overline{p}$ and $\underline{p}$. Symmetry of the problem ($\tilde{p}=p_{0}=1/2$) implies at an optimum $\overline{p}=1-\underline{p}$. The observation follows from the first-order condition $02-\kappa c^{\prime}(p)=$ 0.
While this implies that single-peakedness of $T^{*}(\kappa)$ is not generic, the most common posterior-based information cost functions do have that $c(p)/c^{\prime}(p)$ is quasiconcave on $[1/2,1]$, e.g., entropy costs, $c(p)=p\ln(p)+(1-p)\ln(1-p)-\ln(1/2)$, and quadratic or variance costs, $c(p)=(p-p_{0})^{2}$ — see Figure 3.
@@ 6.  Identifying Complexity
Figure 4 clearly illustrates the result: the higher the indifference point $\tilde{p}$, the more likely the decision-maker is to choose $b$. The effect of problem complexity is then to affect how much more the decision-maker is likely to choose $b$ when the indifference point is increased. It is immediate that the more complex the problem is, insofar as the the choice probability is interior, the stronger the dependence of the choices on the indifference point is, and the steeper the lines are. This is captured by the cross-partial derivative $\partial^{2}\mathbb{P}(\alpha_{\tau}=b)/(\partial\tilde{p}\partial(\mu/\sigma))$ being negative. The proof of this fact follows by obtaining a closed-form characterisation of this derivative using implicit differentiation. The important point is that this result is robust to aggregating over heterogenous agents with different preferences, insofar as these are distributed the same across the two problems.
@@ 7.  Final Remarks
We believe that our results have important implications for the use of response times in the social sciences. While response times provide an important source of information that is able to shed light on the cognitive processes underlying decision-making, our results support existing evidence that, on their own, they are not a reliable indicator of problem complexity or ability. Instead, we propose to use small incentive distortions to infer problem complexity. These are easy to implement in real-life settings and could be used to study which features make specific problems more or less complex, and evaluate if particular policies improve them. In current ongoing work, Gonçalves et al. (2024) not only corroborate the results in Theorem 1 — already evidenced in the above-referenced literature — but also demonstrate that incentive manipulation can indeed be used to reveal problem complexity.
@@ A.5.  Proof of Proposition 1
Suppose that $V_{a}$ has a kink at $p\in(\underline{p},\overline{p})$, i.e., $V_{a}^{\prime}(p^{+})\neq V_{a}^{\prime}(p^{-})$. Then, by convexity, $V_{a}^{\prime}(p^{+})>V_{a}^{\prime}(p^{-})$, and, for any $k\in(V_{a}^{\prime}(p^{-}),V_{a}^{\prime}(p^{+}))$ and significantly high $M>0$, $V_{a}(p^{\prime})\geq V_{a}(p)+k(p^{\prime}-p)+(M/2)(p^{\prime}-p)^{2}=:\psi(p^{\prime})$ for $p^{\prime}$ in a neighborhood of $p$. Then, we would obtain that $\tilde{c}-(\mu/\sigma)^{2}(p(1-p))^{2}M<0$ for sufficiently high $M$, contradicting the fact that $V_{a}$ is a viscosity supersolution. Hence, $V_{a}$ is $\mathcal{C}^{1}$ except possibly at $\{\underline{p},\overline{p}\}$.
Since $V_{a}^{\prime\prime}$ exists a.e., from Crandall et al. (1992, p. 15), the HJB will hold with equality whenever $V_{a}$ is twice differentiable, i.e., almost everywhere on $(\underline{p},\overline{p})$. Furthermore, $0\leq V_{a}^{\prime\prime}\leq(\mu/\sigma)^{-2}\tilde{c}\max\{(\underline{p}(1-\underline{p}))^{-2},(\overline{p}(1-\overline{p}))^{-2}\}$, and so $V_{a}^{\prime\prime}$ is bounded whenever it exists. Take any sequence $(p_{n})\subset(\underline{p},\overline{p})$ such that $p_{n}\to p^{*}\in(\underline{p},\overline{p})$. Then, there is a subsequence $(p_{m})$ such that $V_{a}^{\prime}(p_{m})\to V_{a,\infty}^{\prime}$, $V_{a}^{\prime\prime}(p_{m})\to V_{a,\infty}^{\prime\prime}$, and, by continuity, $0=-\tilde{c}+(\mu/\sigma)^{2}\left(p^{*}(1-p^{*})\right)^{2}V_{a,\infty}^{\prime\prime}$. Hence $(V_{a,\infty}^{\prime},V_{a,\infty}^{\prime\prime})$ belongs to both the super- and subjet of $V_{a}$ at $p^{*}$, implying that $V_{a}$ is twice continuously differentiable everywhere except possibly at $\{\underline{p},\overline{p}\}$.
@@ A.9.  Proof of Proposition 3
Let $\pi_{I}$ denote a fully informative experiment — i.e., assigning probability $p_{0}(\theta)$ to posterior belief $1_{\theta}$ — and $\pi_{U}$ a fully uninformative experiment, assigning probability 1 to the prior $p_{0}$. Take any $\overline{C}>0$ such that $\exists\overline{\kappa}$ such that $\overline{\kappa}C^{*}(\overline{\kappa})>\overline{C}$. Then, there are $\kappa<\overline{\kappa}$ and $\kappa^{\prime}>\overline{\kappa}$ such that $\kappa C^{*}(\kappa)\vee\kappa^{\prime}C^{*}(\kappa^{\prime})<\overline{C}$. To see this note that, by Berge’s maximum theorem, $\kappa C^{*}(\kappa)$ is continuous in $\kappa$. Taking $\kappa_{n}\downarrow 0$, $V(\kappa_{n})\uparrow V(0)=B(\pi_{I})$, and so $V(0)-V(\kappa_{n})=\kappa_{n}C^{*}(\kappa_{n})\downarrow 0$. Take now $\kappa_{n}\uparrow\infty$. Note that $0\leq\lim_{n\to\infty}V(\kappa_{n})/\kappa_{n}=\lim_{n\to\infty}B(\pi^{*}(\kappa_{n}))/\kappa_{n}-C^{*}(\kappa_{n})=\lim_{n\to\infty}-C^{*}(\kappa_{n})\leq 0$, and so it must be that $\pi^{*}(\kappa_{n})\to\pi_{U}$. Consequently, $\kappa_{n}C^{*}(\kappa_{n})\leq B(\pi^{*}(\kappa_{n}))\to 0$, a contradiction. This implies that $\kappa C^{*}(\kappa)$ is non-monotone in $\kappa$ and, therefore, for any strictly increasing $g:\mathbb{R}\to\mathbb{R}$, so is $T^{*}:=g(\kappa C^{*}(\kappa))$. ∎
@@ Appendix B.  Discounting
Furthermore, since payoffs are symmetric (i.e., $u(a,a)-u(b,a)=u(b,b)-u(a,b)$ and $u(a,b)=u(b,a)$), the stopping thresholds will also be symmetric around $1/2$, i.e., $\underline{p}=1-\overline{p}$. Then,